On January 1, 2020, Wildhorse Ltd. had 522,000 common shares outstanding. During 2020, it had the following transactions that affected the common share account:

Feb. 1 Issued 154,000 shares.
Mar. 1 Issued a 20% stock dividend.
May 1 Acquired 174,000 common shares and retired them.
June 1 Issued a 2-for-1 stock split.
Oct. 1 Issued 43,000 shares.


The company’s year end is December 31.

Part 1

Incorrect answer iconYour answer is incorrect.

Determine the weighted average number of shares outstanding as at December 31, 2020. (Round answer to 0 decimal places, e.g. 5,275.)

Weighted average number of shares outstanding =enter the Weighted average number of shares outstanding_______________________

Part 2

Assume that Wildhorse earned net income of $3,553,000 during 2020. In addition, it had 110,000 of 11%, $100 par, non-convertible, non–cumulative preferred shares outstanding for the entire year. Because of liquidity limitations, however, the company did not declare and pay a preferred dividend in 2020.

Calculate earnings per share for 2020, using the weighted average number of shares determined above. (Round answer to 2 decimal places, e.g. 15.25.)

Earnings per share _____________

Part 3

Assume that Wildhorse earned net income of $3,553,000 during 2020. In addition, it had 110,000 of 11%, $100 par, non-convertible, cumulative preferred shares outstanding for the entire year. Because of liquidity limitations, however, the company did not declare and pay a preferred dividend in 2020.

Calculate earnings per share for 2020, using the weighted average number of shares determined above. (Round answer to 2 decimal places, e.g. 15.25.)

Earnings per share $________________

Assume that Wildhorse earned net income of $3,553,000 during 2020. In addition, it had 110,000 of 11%, $100 par, non-convertible, non–cumulative preferred shares outstanding for the entire year. Because of liquidity limitations, however, the company did not declare and pay a preferred dividend in 2020. Assume that net income included a loss from discontinued operations of $421,000, net of applicable income taxes.

Calculate earnings per share for 2020. (Round answers to 2 decimal places, e.g. 15.25.)

Earnings per share

Income from continuing operations

$

Loss from discontinued operations

$

Net income

$
Answer #1

Solution a:

Computation of weighted average outastanding shares
Number of shares Stock Dividend Restatement Stock Split Restatement No. of months Total Months Weighted-average shares
Shares outstanding 522000 1.2 2 1 12 104400
Issued Shares 676000 1.2 2 1 12 135200
Stock Dividend 811200 2 2 12 270400
Reacquired Shares 637200 2 1 12 106200
3-for-1 stock split 1274400 4 12 424800
Reissued shares 1317400 3 12 329350
Total weighted average number of shares outstanding 1370350

Solution b:

Earnings per share = Net income / weighted average number of shares outstanding = $3553000 / 1370350 = $2.59

Solution c:

Preferred Dividend = 110000* 100*11% = $$1,210,000

Earnings per share = (Net income- Preferred dividend) / weighted average number of shares outstanding = ($3553000- $121000) / 1370350 = $1.71

Solution d:

Net Income $35,53,000
Add: Loss from Discontinued Operations $4,21,000
Income from continued operations $39,74,000
Income Statement
For the Year ended Dec 31, 2020
Income from continued operations ($3974000/1370350) $2.90
Loss from Discontinued Operations ($421000/ 1370350) -$0.31
Net Income ($3553000/1370350) $2.59